OpenStudy (anonymous):

Find the volume of the solid formed by rotating the region enclosed by x=0, x=1, y=0, y= 5+x^7, about the y-axis.

OpenStudy (zarkon):

\[\int\limits_{a}^{b}2\pi x f(x)dx\]

OpenStudy (anonymous):

@Zarkon , how do you account for removal of graph below x -axis.

OpenStudy (callisto):

Use shell method... \[Volume = 2\pi \int_{0}^{1} x(5+x^7) dx = 2\pi \int_{0}^{1} (5x+x^8) dx = 2\pi [5/2(x^2) +(1/9)x^9]_{0}^{1} = (47/9)\pi\]

OpenStudy (anonymous):

@Callisto , same question to you as to Zarkon .

OpenStudy (anonymous):

I think you you should divide the final answer by 2

OpenStudy (zarkon):

http://en.wikipedia.org/wiki/Shell_integration

OpenStudy (callisto):

 Length of the shell = f(x) width of the shell = 2pi x thickness of the shell = (delta) x Volume = area x thickness = 2pi int x(f(x) dx